Using the standard molar heat of combustion of hydrogen, methane, and ethane (given below), find the enthalpy change for 2CH4(g) → C2H6(g) + H2(g)
H2 + ½ O2 → H2O ΔHo = -285.8 kJ
CH4 + 2O2 → CO2 + 2H2O ΔHo = -890.4 kJ
C2H6 + (7/2)O2 → 2CO2 + 3H2O ΔHo = -1559.9 kJ
Answer #1

Given reactions are :

H2 + ½ O2 → H2O ΔHo = -285.8 kJ......(1)

CH4 + 2O2 → CO2 + 2H2O ΔHo = -890.4 kJ . .....]*2

2CH4 + 4O2 → 2CO2 + 4H2O ΔHo = -1780.8 kJ.....(2)

C2H6 + (7/2)O2 → 2CO2 + 3H2O ΔHo = -1559.9 kJ....(3)

Subtracting (1) and (3) from (2) , we get

2CH4 + 4O2 - H2 - ½ O2 - C2H6 - (7/2)O2 → 2CO2 + 4H2O - 2CO2 - 3H2O - H2O   ΔHo = -1780.8 kJ +285.8 + 1559.9 kJ

2CH4 → H2 + C2H6 ΔHo = 64.9 kJ